if my equation is CH4 + 2O2 = CO2 + 2H2O and i have 1.246 moles of oxygen, how would i use this information to find the mass of another substance in the equation?
1
Expert's answer
2020-06-04T10:38:29-0400

For example, according to the equation 1 mol of CH4 reacts with 2 mol of O2. Then 1.246 mol of O2 reacts with

n(CH4) = n(O2)/2 = 1.246/2 = 0.623 mol

Mass of CH4:

m(CH4) = n*M = 0.623mol*16g/mol = 9.968 g